. Find the value of x and the value of y in the following equation. 𝟏 𝒙+π’šπ’Š - 𝟏 𝟏+π’Š = 2 – 3i


1
Expert's answer
2021-09-08T15:23:02-0400

1x+yi-11+I=2-3i

Collecting like terms,

1x-11=2

1x=2+11

1x=13

Hence, x=13


yi+i=-3i

yi=-3i-i

yi=-4i

Hence, y=-4